OpenStudy (anonymous):

What is the equation of the graphed line? Hint: Determine the slope of the line. y = x Graph of a line passing through the origin and Quadrants II and IV. Line passes through (1, -5).

OpenStudy (texaschic101):

lets find the slope by using the slope formula and 2 sets of points. slope(m) = (y2 - y1) / (x2 - x1) (0,0)...x1 = 0 and y1 = 0 (1,-5)...x2 = 1 and y2 = -5 now we sub slope(m) = (-5 - 0) / (1 - 0) slope(m) = -5/1 or just -5 now we use y = mx + b slope(m) = -5 use either of your points...(1,-5)...x = 1 and y = -5 now we sub and find b, the y intercept -5 = -5(1) + b -5 = -5 + b -5 + 5 = b 0 = b so the equation of the line is : y = -5x + 0...or just y = -5x